OpenStudy (anonymous):

2ln x = ln(2x-1) log4 (x-2)=2

OpenStudy (anonymous):

2ln x = ln(2x-1) ln (x^2) = ln(2x-1) x^2 = e^(ln(2x-1)) x^2 = 2x-1 x^2-2x+1 = 0 (x-1)^2 = 0 x = 1